OpenStudy (anonymous):

Barbara opens a credit card with an APR of 19.99%. How much is charged in interest this month if her balance is $1250? $19.99 $20.82 $249.88 $16.66

OpenStudy (amistre64):

1250 = old balance (1+r/12) interest added is the difference between the old and new balance

OpenStudy (amistre64):

interest added = 1250 - old balance old balance = 1250/k i = 1250 - 1250/k i = 1250( 1 - 1/k)

OpenStudy (amistre64):

or if im misreading the question, the old balance is 1250 and they want to know how much interest is added to get the new balance ... or just 1250 - 1250k

OpenStudy (amistre64):

either way i get about 20 and change